Advertisements
Advertisements
Question
A and B are two events such that P (A) = 0.25 and P (B) = 0.50. The probability of both happening together is 0.14. The probability of both A and B not happening is
Options
0.39
0.25
0.11
none of these
Advertisements
Solution
0.39
P (A) = 0.25 and P (B) = 0.50
P(A \[\cap\] B) = 0.14
∴ Required probability = 1 - P(A∪B)
= 1 - [P(A) + P(B) - P(A \[\cap\]B)]
= 1 - [0.25 + 0.50 - 0.14]
= 1 - 0.61 = 0.39
APPEARS IN
RELATED QUESTIONS
If E and F are events such that P(E) = `1/4`, P(F) = `1/2` and P(E and F) = `1/8`, find
- P(E or F)
- P(not E and not F).
A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(not A).
In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.
A die has two faces each with number ‘1’, three faces each with number ‘2’ and one face with number ‘3’. If die is rolled once, determine
- P(2)
- P(1 or 3)
- P(not 3)
Three letters are dictated to three persons and an envelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.
A dice is thrown. Find the probability of getting a prime number
In a simultaneous throw of a pair of dice, find the probability of getting:
8 as the sum
In a simultaneous throw of a pair of dice, find the probability of getting a doublet
In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers
In a simultaneous throw of a pair of dice, find the probability of getting a sum greater than 9
In a simultaneous throw of a pair of dice, find the probability of getting a sum less than 7
In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 7
In a simultaneous throw of a pair of dice, find the probability of getting odd number on the first and 6 on the second
In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.
In a single throw of three dice, find the probability of getting a total of 17 or 18.
Three coins are tossed together. Find the probability of getting at least two heads
Two dice are thrown. Find the odds in favour of getting the sum 5.
Two dice are thrown. Find the odds in favour of getting the sum What are the odds against getting the sum 6?
A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn at random. From the box, what is the probability that all are blue?
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white and odd numbered .
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| \[\frac{1}{3}\] | \[\frac{1}{5}\] | \[\frac{1}{15}\] | ...... |
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| 0.35 | .... | 0.25 | 0.6 |
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| 0.5 | 0.35 | ..... | 0.7 |
If A and B are two events associated with a random experiment such that
P (A ∪ B) = 0.8, P (A ∩ B) = 0.3 and P \[(\bar{A} )\]= 0.5, find P(B).
There are three events A, B, C one of which must and only one can happen, the odds are 8 to 3 against A, 5 to 2 against B, find the odds against C
One of the two events must happen. Given that the chance of one is two-third of the other, find the odds in favour of the other.
In a single throw of two dice, find the probability that neither a doublet nor a total of 9 will appear.
The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English Examination is 0.75. What is the probability of passing the Hindi Examination?
One of the two events must occur. If the chance of one is 2/3 of the other, then odds in favour of the other are
If the probability of A to fail in an examination is \[\frac{1}{5}\] and that of B is \[\frac{3}{10}\] . Then, the probability that either A or B fails is
A box contains 10 good articles and 6 defective articles. One item is drawn at random. The probability that it is either good or has a defect, is
Two dice are thrown simultaneously. The probability of getting a pair of aces is
An urn contains 9 balls two of which are red, three blue and four black. Three balls are drawn at random. The probability that they are of the same colour is
Five persons entered the lift cabin on the ground floor of an 8 floor house. Suppose that each of them independently and with equal probability can leave the cabin at any floor beginning with the first, then the probability of all 5 persons leaving at different floor is
